a sum of rs 2000 is split between two investments, one paying 7% interest and the other paying 8%. if the return on 8% investment exceeds that on the 7% investment by rs 40, how much is invested at each rate?
Answer by mananth(12270)

Part I 8.00% per annum ------------- x
Part II 7.00% per annum ------------ y
total investment 2000
Interest----- from I = 8.00% x
Interest from II= 7.00% y
greater by 40
8.00% x = 7.00% y + 40
multiply by 100
8 x - -7 y = 4000 --------1
x + y = 2000 -------- 2
eliminate y
Multiply 1 by 1
Multiply 2 by 7
we get
8 x - -7 y = 4000
7 x - 7 y = 14000
15 x = 18000
/ 15
x= 1200
Plug the value of x in in 1
9600 -7 y = 4000
-7 y = 4000 -9600
-7 y = -5600
/ -7
y= 800
Rs 1200 at 8.00%
Rs 800 at 7.00%

CHECK
96 = 56 + 40
m.ananth@hotmail.ca